OpenStudy (anonymous):

simplify- sqrt24-3sqrt+2sqrt20

OpenStudy (anonymous):

\[\sqrt{24}-3\sqrt{?}+2\sqrt{20}\] is something missing? \[2\sqrt{6}-3\sqrt{?}+4\sqrt{5}\]

OpenStudy (anonymous):

sqrt of 45 sorry

OpenStudy (anonymous):

\[2\sqrt{6}-9\sqrt{5}+4\sqrt{5}\] \[2\sqrt{6}-5\sqrt{5}\]